OpenStudy (anonymous):

A fishing entushiast buys fifty maggots and twenty worms for $1.10, and her mother buys thirty maggots and forty worms for $1.50 Find the cost of one maggot and one worm maggot = x worm = y A fishing entushiast buys fifty maggots and twenty worms for $1.10, and her mother buys thirty maggots and forty worms for $1.50 Find the cost of one maggot and one worm maggot = x worm = y @Mathematics

OpenStudy (anonymous):

i just wanted it turned to a simultaneous equation as i will try to solve it myself

OpenStudy (anonymous):

Let maggot' price be x Let worm' price be y 50x + 20y =1.1 30x + 40y = 1.5

OpenStudy (anonymous):

100x + 40y = 2.2 - 30x + 40y = 1.5 70x = 0.7 x = 0.01 y = 0.03